1. If 4.000 grams of hydrogen peroxide is placed within a sealed 250 milliliter container at 500 K.

What is the pressure of the oxygen gas produced in atmospheres?

2
1
Expert's answer
2020-09-15T13:50:22-0400

Molar mass of "H_2O_2= 34 \ g"

Moles of "H_2O_2= \frac{4}{34}"

"2H_2O_2\to 2H_2O + O_2"

So ,moles of oxygen produced will be half of hydrogen peroxide .

"n= \frac{2}{34}"

"T=500K,V= 250\ ml = 0.25 \ L"

Using ideal gas equation,

"PV= nRT"

"P\times 0.25= \frac{2}{34}\times 0.0821 \times 500\implies"

"P=9.66 \ atm"